OpenStudy (anonymous):

Mr Frankel bought 5 tickets to a pupet show and spent 26.00. He bought some child tickets for 2.00 each and some adult tickets for 10.00 each. Write a system of equations to complete this problem

OpenStudy (lilg132):

TP = a(x) + c(y) 26 = a(2) + c(3)